OpenStudy (anonymous):

Find the present value of $10,000 if the interest is paid at a rate of 9% per year, compounded semiannually for 3 years.

OpenStudy (anonymous):

compounded semi-annually the formula is: \(\large A=P(1+\frac{rate}{2})^{2*t} \) so, \(\large A=10000(1+\frac{0.09}{2})^{2*(3)} \)